BELLA38:

Arlene is comparing the prices of two truck rental companies. Company A charges $7 per hour and an additional $30 as service charges. Company B charges $8 per hour and an additional $40 as service charges. Part A: Write an equation to represent each company's total charges for renting a truck for a certain number of hours. For both equations (one for Company A and one for Company B), define the variable used. (4 points) Part B: Which company would charge less for renting a truck for 8 hours? Justify your answer. (3 points) Part C: How much money is saved by using the services of Company A instead of Company B to rent a truck for 2 hours?

TheSmartOne:

Hello @BELLA38

TheSmartOne:

How would you write the equations? Remember the form y = mx + b m is the slope, it's the rise/run It's the number we're constantly changing for every single increase in x b is our y-intercept it's the constant

TheSmartOne:

So for your question, we are trying to write an equation for the price of a rental truck if we say 'm' is how much it costs per hour and 'y' is our total cost after 'x' number of hours then can you figure out how we would write the equation? Don't forget to include the constant, the base starting price (our 'b' value)

TheSmartOne:

what would the equations be?

TheSmartOne:

That's part A. For part B: Once you have your equations, they're going to be in that form of y = mx + b And you just have to put in the values they gave you in the question as the numbers for m and b, to get your equations. To answer part b, you just plug in 8 into x and solve for y. You do that for both equations and you can find out which company will charge less for renting a truck for 8 hours.

TheSmartOne:

Part 3 is similar where instead of plugging in 8, you plug in 2 But now you also have to find how much money you're saving So you just do Higher cost - lower cost And that's how much money you're saving by going with the company that charges less

TheSmartOne:

Part C** not part 3 but I'm sure you get the idea :)

TheSmartOne:

Anyways, let me know how far you get!

BELLA38:

THANKYOU!

TheSmartOne:

My pleasure!! Also welcome to QuestionCove!! If you want me to check your answers or want me to clarify something more, do let me know!

BELLA38:

Will do , I really appreciate it !

BELLA38:

@TheSmartOne 35 =7x + 30 is this correct for part A or do i have to put the 5 instead of the x?

BELLA38:

104= 8x + 40 and is this correct for part b of part A?

BELLA38:

or should i put the 8 instead of the x

TheSmartOne:

I'm not sure where you got the 35 and 104 It should just be y = And then it'll be correct

TheSmartOne:

And that would be the equations for company A and company B respectively

BELLA38:

Great thankyou !!

BELLA38:

Part B : Company A would charge less, because they charge 56 for 8 hours plus the service fee of 30$making it 86 $ total & Company B charges 104$ total for 8 hours plus the 40 $ service fee.. Company A: 7x8 =56+30=86 Company B: 8x 8 =64 +40= 104

BELLA38:

is that correct for part b?

BELLA38:

Part C : Company A: 7 x 2= 14 +30 = 44$ for comapny A Company B: 8 x 2= 16 + 40 =56 for company B So we would be saving 12 $ by going with Company A instead of Company B when renting the truck for 2 hours.

TheSmartOne:

Yup!! Good job!

BELLA38:

Is this correct ? I believe so just want to make sure .. Again thank you so much for your time & help !!

TheSmartOne:

Remember for part b to clarify which company is cheaper And for part A to mention what x means and what y means (I told you what this one was in one of my first responses)

TheSmartOne:

And of course! Anytime! :)

BELLA38:

Great Thank Youu !!! couldn't of done it without your help !!

BELLA38:

Great thankyou couldnt of done it without your help!!

TheSmartOne:

Of course! It was my pleasure :)